6. Let R be a non-trivial ring with unity. Suppose x E R satisfies x2 = 0. Show that x – 1 and x +1 are units in R.
Answer #1

6. let R be a non- trivial ring with unity let XERR such that x²=0. Now, (1+x), (1-x) E R (1+x) (1-x) = 1- x2 = 1-0 =1. = (1+